What is the molarity of a solution prepared by diluting 10.0mL of 2.7 M NaOH to 60.0 mL?
1
Expert's answer
2019-01-10T04:05:27-0500

n = 2.7*0.01 = 0.027 moles

C = 0.027/0.06 = 0.45M

Answer. 0.45M